OpenStudy (anonymous):

if the 5th term of a GP is 2 then the product of its first 9 terms is?

OpenStudy (anonymous):

@satellite73 @mukushla

OpenStudy (anonymous):

@amistre64

OpenStudy (amistre64):

your missing an element, or are you spose to keep it generic?

OpenStudy (anonymous):

My question is correct the answer should be 512

OpenStudy (amistre64):

(a1) * (a1*r) * (a1*r*r ) * (a1*r*r*r) * (a1*r*r*r*r) ... =2

OpenStudy (callisto):

I'm thinking if it works... 5th term of a GP is 2 => \(ar^4 = 2\) => \(a= \frac{2}{r^4}\) General term of a GP = \(ar^{n-1}\) product of its first 9 terms = \((ar^{1-1})\times (ar^{2-1}) \times ...\times (ar^{9-1})\) = \(a^9r^{0+1+2+...+8}\) = \((\frac{2}{r^4})^9 r^{36}\) = \(2^9\)

OpenStudy (amistre64):

callistos is fancier :)

OpenStudy (anonymous):

thxxx

OpenStudy (callisto):

@Yahoo! Welcome.... Do you understand my workings???

OpenStudy (anonymous):

yup!!

OpenStudy (callisto):

Glad to hear that :)

OpenStudy (anonymous):

@Callisto doubt

OpenStudy (callisto):

Yes?

OpenStudy (anonymous):

hw r^34?

OpenStudy (callisto):

It's r^(36) Since it's r^(0+1+2+..+8) 0+1+2+...+8 = (1+8)x8 / 2 = 36 So, r^(0+1+2+..+8) = r^36

OpenStudy (anonymous):

got it thxx

OpenStudy (callisto):

Welcome!!~